How much of an 8% saline solution should be added to 600 ml of a 3% solution to produce a 5% solution?
Answer by edjones(8007)

.08 * x =.08x
.03*600 = 18
.05*(x+600)=.05x+30
.
.08x+18=.05x+30
.03x=12
x=400 ml
.
Ed